# FreezeAsGuard: Mitigating Illegal Adaptation of Diffusion Models via Selective Tensor Freezing

### v1

@@ 3.1 Mask Learning in the Upper-level Loop
To solve the upper-level multi-objective optimization in Eq. (2), we adopt linear scalarization [33] to convert it into a single objective $\mathcal{L}_{upper}$ via a weighted summation with weights $(\lambda_{1},\lambda_{2})$:
$\mathcal{L}_{upper}=-\lambda_{1}\mathcal{L}_{\boldsymbol{\theta}^{*}(\mathbf{m})}(\mathbf{x}_{illegal})+\lambda_{2}\mathcal{L}_{\boldsymbol{\theta}^{*}(\mathbf{m})}(\mathbf{x}_{innocent}),\vspace{-0.05in}$ (5)
to involve training samples in both illegal and innocent domains when learning $\mathbf{m}$. Intuitively, $(\lambda_{1},\lambda_{2})$ should ensure that gradient-based feedbacks from the two loss terms are not biased by any inequality between the amounts of $\mathbf{x}_{illegal}$ and $\mathbf{x}_{innocent}$. In practice, we can set $(\lambda_{1},\lambda_{2})=(1,1)$ because computing expectation in the diffusion loss in Eq. (1) has normalized the loss to sample-wise magnitude.
Besides, $\mathbf{x}_{illegal}$ and $\mathbf{x}_{innocent}$ could contain some knowledge in common, and masked learning from such data may hence affect model adaptation in innocent domains. To address this problem, we add a sparsity constraint $\mathcal{L}_{sparsity}$ to $\mathcal{L}_{upper}$ to better control of the mask’s mitigation power:
$\mathcal{L}_{sparsity}=\|\mathbf{1}^{\top}\mathbf{m}/N-\rho\|_{2}^{2},\vspace{-0.05in}$ (6)
where $N$ is the total number of tensors and $\mathbf{1}^{\top}\mathbf{m}/N$ measures the proportion of tensors being frozen. By minimizing $\mathcal{L}_{sparsity}$, the achieved ratio of tensor freezing should approach the given $\rho$ specified by model publisher. In this way, we can apply gradient descent to minimize $\mathcal{L}_{upper}$ and iteratively refine $\mathbf{m}$ towards optimum, and we will experimentally investigate the optimal value of $\rho$ in Section 4.
@@ 3.2 Model Fine-tuning in the Lower-level Loop
Effectiveness of mask learning at the upper level relies on timely feedback from the lower-level fine-tuning. Every time the mask has been updated by an iteration in the upper level, the lower-level loop should adopt the updated mask into fine-tuning, and return the fine-tuned model tensors and the correspondingly updated loss value as feedback to the upper level. Similar to Eq. (5), the fine-tuning objective is the summation of diffusion losses for illegal and innocent domains:
$\mathcal{L}_{lower}=\mathcal{L}_{\boldsymbol{\theta}^{*}(\mathbf{m})}(\mathbf{x}_{illegal})+\mathcal{L}_{\boldsymbol{\theta}^{*}(\mathbf{m})}(\mathbf{x}_{innocent}).\vspace{-0.05in}$ (7)
3.3 Improving the Compute Efficiency of Bilevel Optimization
Solving bilevel optimization has known to be computationally expensive, due to the repeated switches between the upper-level and lower-level loops [46; 59]. Rigorously, as shown in Figure 6 - Left, every time when the mask has been updated, the model should be fine-tuned with a sufficient number of iterations until convergence, before the next update of the mask. In this way, the mask can be optimized to maximize the loss value when fine-tuning converges. However, in practice, doing so is extremely expensive, given the possibly large number of mask updates before reaching $\mathbf{m}^{*}$.
@@ 3.3 Improving the Compute Efficiency of Bilevel Optimization
Instead, as shown in Figure 6 - Right, we observe that the fine-tuning loss typically drops fast in the first few iterations and then starts to violently fluctuate (see details in Appendix B). Hence, every time in the lower-level loop of model fine-tuning, we do not wait for the fine-tuning loss to converge, but instead only fine-tune the model for the first a few iterations before updating the mask to the upper-level loop of mask learning. After the model update, the fine-tuned model weights are inherited to the next loop of model fine-tuning, to ensure consistency and improve convergence. In this way, the optimization only needs one fine-tuning process, during which the mask can be updated with shorter intervals but higher learning quality. Details of deciding such a number of fine-tuning iterations are in Appendix B.
Further, to perform bilevel optimizations, three versions of diffusion model weights, i.e., $\boldsymbol{\theta}(\mathbf{m})$, $\boldsymbol{\theta}_{pre}$ and $\boldsymbol{\theta}_{ft}$, will be maintained for gradient computation as shown in Figure 5. This could significantly increase the memory cost due to large sizes of today’s diffusion models. To reduce such memory cost, we instead maintain only two versions of model weights, namely $\boldsymbol{\theta}(\mathbf{m})$ and $\boldsymbol{\theta}_{d}=\boldsymbol{\theta}_{pre}-\boldsymbol{\theta}_{ft}$. According to Eq. (12) and (11) in Appendix A, the involvement of both $\boldsymbol{\theta}_{pre}$ and $\boldsymbol{\theta}_{ft}$ can be removed by plugging $\boldsymbol{\theta}_{d}$ into the gradient descent calculation. More specifically, for a given model tensor $i$, the gradient descent to update the corresponding mask $m_{i}$ in the upper-level optimization is:
$w_{i}\leftarrow w_{i}-\eta_{1}\left<\frac{\partial{\mathcal{L}_{upper}}}{\partial{\theta(m)_{i}}},\theta_{d}^{(i)}\right>\frac{1}{T}\sigma\left(\frac{w_{i}}{T}\right)\sigma\left(1-\frac{w_{i}}{T}\right),\ \ \ \ m_{i}\leftarrow,\sigma\left(\frac{w_{i}}{T}\right)$ (8)
where $\eta_{1}$ controls the step size of updates. On the other hand, computing the update of $\boldsymbol{\theta}(\mathbf{m})$ and $\boldsymbol{\theta}_{d}$ at the lower level should apply the chain rule over Eq. (12):
@@
Algorithm 1 Optimizing Freezing Strategy in FreezeAsGuard
1:Illegal and innocent domain data $(\mathcal{D}_{illegal},\mathcal{D}_{innocent})$, step size $\eta_{1}$ and $\eta_{2}$, model weights $\boldsymbol{\theta}_{pre}$ and $\boldsymbol{\theta}_{ft}$
@@
5:         $(\mathbf{x}_{illegal},\mathbf{x}_{innocent})\leftarrow$ RandomSample$(\mathcal{D}_{illegal},\mathcal{D}_{innocent})$
6:         $\frac{\partial{\mathcal{L}_{lower}}}{\partial{\boldsymbol{\theta}(\mathbf{m})}}\leftarrow$ Backprop$(\mathbf{x}_{illegal},\mathbf{x}_{innocent},\mathcal{L}_{lower},\boldsymbol{\theta}(\mathbf{m}))$
7:         $(\boldsymbol{\theta}_{d},\boldsymbol{\theta}(\mathbf{m}))\leftarrow$ ComputeUpdate$\left(\frac{\partial{\mathcal{L}_{lower}}}{\partial{\boldsymbol{\theta}(\mathbf{m})}},\mathbf{m},\boldsymbol{\theta}_{d},\boldsymbol{\theta}(\mathbf{m}),\eta_{1}\right)$     // Refer to Eq. (9) and (10)
@@
9:    $(\mathbf{x}_{illegal},\mathbf{x}_{innocent})\leftarrow$ RandomSample$(\mathcal{D}_{illegal},\mathcal{D}_{innocent})$
10:    $\frac{\partial{\mathcal{L}_{upper}}}{\partial{\boldsymbol{\theta}(\mathbf{m})}}\leftarrow$ Backprop$(\mathbf{x}_{illegal},\mathbf{x}_{innocent},\mathcal{L}_{upper},\boldsymbol{\theta}(\mathbf{m}))$
11:    $\mathbf{m}\leftarrow$ ComputeUpdate$\left(\frac{\partial{\mathcal{L}_{upper}}}{\partial{\boldsymbol{\theta}(\mathbf{m})}},\mathbf{m},\boldsymbol{\theta}_{d},\eta_{2}\right)$     // Refer to Eq. (8)
@@ 4 Experiments
Models & Datasets. In our evaluation, we use three popular open-source diffusion models: SD v1.4 [7], v1.5 [8] and v2.1 [9], all with 1B parameters. These diffusion models have architecture variations and are pre-trained on the LAION-5B dataset [48] with different training configurations.
We focus on a challenging mitigation of illegal model adaptation that has high social impact: generating fake portraits of public figures [24; 26]. We use a self-collected portrait image dataset with synthetic text prompts (FF25) as the illegal domain, and use two other public datasets of logo images (Logo) and fashion products (Clothes) as innocent domains [^3] Most existing work uses toy image datasets (e.g., CIFAR [36] and subsets of ImageNet [47]) with non-living objects, anonymous human faces [41] and short text descriptions, which are far from practical model adaptations.. Details of these datasets are as follows:
  • Famous-Figures-25 (FF25): We use AutoCrawler [2] to collect 8,703 publicly available portrait images of 25 public figures from the Web, with 400-1,300 images for each figure. For each image, we use a pre-trained BLIP2 image captioning model [38] to generate its text description, by using the prompt “a photo of <person_name> which shows” to avoid hallucination. More details are described in Appendix C.
  • Modern-Logo-v4 [4] (Logo): It contains 803 logo images that are labeled with informative text descriptions. The logos are minimalist, meeting modern design requirements and reflecting the corresponding company’s industry. Some sample images are in Appendix E.
  • H&M-Clothes [3] (Clothes): It contains 1k images of fashion products from H&M Group with informative text descriptions, including color, texture, and functionality of the product to help customers’ decision making. Some sample images are in Appendix E.
Baseline Schemes. We compare FreezeAsGuard with the following baselines, including full fine-tuning, random freezing, and two competitive unlearning schemes. Existing data poisoning methods [54; 57; 49] are not applicable because all data used in our evaluations are publicly available online.
@@
  • UCE [25]: It applies unlearning to guide the learned knowledge about illegal domains to irrelevant or more generic. In this way, it reduces the model’s fine-tuning power in illegal domains by removing related knowledge in the pre-trained model.
@@
  • IMMA [59]: It reinitializes the model weights so that it is hard for users to conduct effective fine-tuning on the reinitialized model, in both illegal and innocent domains.
Evaluation Setup. Details about configurations of training and testing data for mask learning and model fine-tuning, as well as hyperparameter setup, are in Appendix F. We use multiple image quality metrics, including FID [30], CLIP score [29], and TOPIQ-FR [15], to evaluate the quality of generated images. For each subject (public figure) in the illegal domain, we use FID to measure the distance between the distributions of training images and generated images. For innocent domains, we use the CLIP score to measure the goodness of alignment between generated images and the prompt. For each prompt, the results are averaged from 100 randomly generated images.
Method Illegal (10 subjects) Subject: Angela Merkel Subject: Bill Gates Innocent (Logo)
FID ($\downarrow$) TOPIQ ($\uparrow$) FID ($\downarrow$) TOPIQ ($\uparrow$) FID ($\downarrow$) TOPIQ ($\uparrow$) CLIP ($\uparrow$) TOPIQ ($\uparrow$)
Ground Truth - - - - - - 32.24 -
Full FT 143.58 0.097 165.99 0.070 135.43 0.089 32.83 0.097
UCE 151.53 0.097 170.04 0.085 137.19 0.088 30.27 0.088
IMMA 148.05 0.095 167.68 0.077 137.71 0.089 28.71 0.094
Random-1% 152.42 0.097 172.84 0.083 139.55 0.090 30.44 0.107
FreezeAsGuard-1% 153.11 0.095 178.75 0.077 147.73 0.088 30.94 0.118
Random-5% 149.62 0.097 168.22 0.084 134.84 0.088 30.91 0.100
FreezeAsGuard-5% 151.82 0.097 175.63 0.088 136.86 0.085 29.65 0.088
Random-10% 149.10 0.097 168.50 0.085 147.75 0.088 30.88 0.105
FreezeAsGuard-10% 154.51 0.097 170.85 0.081 146.84 0.090 30.04 0.101
Random-20% 152.65 0.097 166.79 0.083 140.95 0.090 29.80 0.096
FreezeAsGuard-20% 156.09 0.095 175.41 0.086 154.33 0.083 30.26 0.097
Random-30% 152.21 0.098 166.26 0.086 124.99 0.096 28.92 0.102
FreezeAsGuard-30% 155.98 0.098 175.77 0.086 150.10 0.091 30.47 0.102
Random-40% 156.08 0.097 172.48 0.084 134.63 0.090 30.03 0.092
FreezeAsGuard-40% 156.02 0.097 169.43 0.086 138.10 0.090 30.91 0.105
Random-80% 157.56 0.094 181.49 0.086 147.22 0.093 28.42 0.101
FreezeAsGuard-80% 156.89 0.094 179.32 0.086 149.89 0.090 29.67 0.104
Table 2: Quantitative results on illegal (10 public figure subjects in FF25 dataset) and innocent domains (LoGo dataset), where worse image quality scores indicate stronger mitigation power in the domain. Results over two specific subjects, Angela Merkel and Bill Gates, are also shown as examples.
4.1 Quantitative Results of Mitigation Power in Illegal and Innocent Domains
To evaluate the basic performance of FreezeAsGuard, we construct the illegal domain as data samples from 10 random subjects in FF25 dataset, and use the Logo dataset as innocent domain. We follow the method in Appendix F to split these data for mask learning and model fine-tuning, and use SD v1.5 with different freezing ratios ($\rho$%) of model tensors, denoting as (FreezeAsGuard-$\rho$%).
As shown in Table 2, FreezeAsGuard can effectively mitigate model adaptation in the illegal domain, measured by the quality of images generated by the fine-tuned model, by a large margin (14% in FID) compared to Full FT [^4] TOPIQ doesn’t show large variations even on baselines, because it only measures standard distortions (e.g., Gaussian blur and lossy compression) but not distortions due to insufficient model representation power.. When $\rho$ varies from 1% to 80%, with the optimally set of frozen tensors, FreezeAsGuard always outperforms baselines of unlearning schemes (UCE [25] and IMMA [59]), which cannot prevent relearning knowledge in illegal domains with new training data. FreezeAsGuard also maintains better quality of model adaptation in the innocent domain, compared to baselines.

### v2

@@ 3.1 Mask Learning in the Upper-level Loop
To solve the upper-level optimization in Eq. (2), we adopt linear scalarization [29] to convert it into a single objective $\mathcal{L}_{upper}$ via a weighted summation with weights $(\lambda_{1},\lambda_{2})$:
$\mathcal{L}_{upper}=-\lambda_{1}\mathcal{L}_{\boldsymbol{\theta}^{*}(\mathbf{m})}(\mathbf{x}_{illegal})+\lambda_{2}\mathcal{L}_{\boldsymbol{\theta}^{*}(\mathbf{m})}(\mathbf{x}_{legal}),$ (5)
to involve training samples in both illegal and legal classes when learning $\mathbf{m}$. $(\lambda_{1},\lambda_{2})$ should ensure that gradient-based feedbacks from the two loss terms are not biased by inequality between the amounts of $\mathbf{x}_{illegal}$ and $\mathbf{x}_{legal}$, and their values should be proportionally set based on these amounts.
Besides, $\mathbf{x}_{illegal}$ and $\mathbf{x}_{legal}$ could contain some knowledge in common, and masked learning from such data may hence affect model adaptation in legal classes. To address this problem, we add a sparsity constraint $\mathcal{L}_{sparsity}$ to $\mathcal{L}_{upper}$ to better control of the mask’s mitigation power:
$\mathcal{L}_{sparsity}=\|\mathbf{1}^{\top}\mathbf{m}/N-\rho\|_{2}^{2},$ (6)
where $N$ is the number of tensors and $\mathbf{1}^{\top}\mathbf{m}/N$ measures the proportion of tensors being frozen. By minimizing $\mathcal{L}_{sparsity}$, the achieved ratio of tensor freezing should approach the given $\rho$. In this way, we can apply gradient descent to minimize $\mathcal{L}_{upper}$ and iteratively refine $\mathbf{m}$ towards optimum.
@@ 3.2 Model Fine-tuning in the Lower-level Loop
Effectiveness of mask learning at the upper level relies on timely feedback from the lower-level fine-tuning. Every time the mask has been updated by an iteration in the upper level, the lower-level loop should adopt the updated mask into fine-tuning, and return the fine-tuned model tensors and the correspondingly updated loss value as feedback to the upper level. Similar to Eq. (5), the fine-tuning objective is the summation of diffusion losses for illegal and legal domains:
$\mathcal{L}_{lower}=\mathcal{L}_{\boldsymbol{\theta}^{*}(\mathbf{m})}(\mathbf{x}_{illegal})+\mathcal{L}_{\boldsymbol{\theta}^{*}(\mathbf{m})}(\mathbf{x}_{legal}).$ (7)
@@
3.3 Towards Efficient Bilevel Optimization
Solving bilevel optimization is computationally expensive, due to the repeated switches between upper-level and lower-level loops [47; 65]. Rigorously, as shown in Figure 6 - Left, every time when the mask has been updated, the model should be fine-tuned with a sufficient number of iterations until convergence, before the next update of the mask. However, in practice, doing so is extremely expensive.
Instead, as shown in Figure 6 - Right, we observe that the fine-tuning loss typically drops fast in the first few iterations and then violently fluctuates (see Appendix B). Hence, every time in the lower-level loop of model fine-tuning, we do not wait for the loss to converge, but only fine-tune the model for the first few iterations before updating the mask to the upper-level loop of mask learning. After the model update, the fine-tuned model weights are inherited to the next loop of model fine-tuning, to ensure consistency and improve convergence. Hence, the optimization only needs one fine-tuning process, during which the mask can be updated with shorter intervals but higher learning quality. Details of deciding such a number of iterations are in Appendix B.
Further, to perform bilevel optimizations, three versions of diffusion model weights, i.e., $\boldsymbol{\theta}(\mathbf{m})$, $\boldsymbol{\theta}_{pre}$ and $\boldsymbol{\theta}_{ft}$, will be maintained for gradient computation. This could significantly increase the memory cost due to large sizes of diffusion models. To reduce such memory cost, we instead maintain only two versions of model weights, namely $\boldsymbol{\theta}(\mathbf{m})$ and $\boldsymbol{\theta}_{d}=\boldsymbol{\theta}_{pre}-\boldsymbol{\theta}_{ft}$. According to Appendix A, the involvement of both $\boldsymbol{\theta}_{pre}$ and $\boldsymbol{\theta}_{ft}$ can be removed by plugging $\boldsymbol{\theta}_{d}$ into the gradient descent calculation. More specifically, for a given model tensor $i$, the gradient descent to update the corresponding mask $m_{i}$ in the upper-level optimization is:
$w_{i}\leftarrow w_{i}-\eta_{1}\left<\frac{\partial{\mathcal{L}_{upper}}}{\partial{\theta(m)_{i}}},\theta_{d}^{(i)}\right>\frac{1}{T}\sigma\left(\frac{w_{i}}{T}\right)\sigma\left(1-\frac{w_{i}}{T}\right),$ (8)
where $\eta_{1}$ controls the step size of updates and $m_{i}$ is updated as $\sigma\cdot w_{i}/{T}$. Further, computing the update of $\boldsymbol{\theta}(\mathbf{m})$ and $\boldsymbol{\theta}_{d}$ at the lower level should apply the chain rule:
@@ 3.3 Towards Efficient Bilevel Optimization
Algorithm 1 Freezing Strategy in FreezeAsGuard
1:Illegal and legal class data $(\mathcal{C}_{illegal},\mathcal{C}_{legal})$, step size $\eta_{1}$ and $\eta_{2}$, model weights $\boldsymbol{\theta}_{pre}$ and $\boldsymbol{\theta}_{ft}$
@@
5:         $(\mathbf{x}_{illegal},\mathbf{x}_{legal})\leftarrow$ Sample$(\mathcal{C}_{illegal},\mathcal{C}_{legal})$
6:         $\frac{\partial{\mathcal{L}_{lower}}}{\partial{\boldsymbol{\theta}(\mathbf{m})}}\leftarrow$ Backprop$(\mathbf{x}_{illegal},\mathbf{x}_{legal},\mathcal{L}_{lower},\boldsymbol{\theta}(\mathbf{m}))$
7:         $(\boldsymbol{\theta}_{d},\boldsymbol{\theta}(\mathbf{m}))\leftarrow$ Update$\left(\frac{\partial{\mathcal{L}_{lower}}}{\partial{\boldsymbol{\theta}(\mathbf{m})}},\mathbf{m},\boldsymbol{\theta}_{d},\boldsymbol{\theta}(\mathbf{m})\right)$     // Refer to Eq. (9) and (10)
@@
9:    $(\mathbf{x}_{illegal},\mathbf{x}_{legal})\leftarrow$ Sample$(\mathcal{C}_{illegal},\mathcal{C}_{legal})$
10:    $\frac{\partial{\mathcal{L}_{upper}}}{\partial{\boldsymbol{\theta}(\mathbf{m})}}\leftarrow$ Backprop$(\mathbf{x}_{illegal},\mathbf{x}_{legal},\mathcal{L}_{upper},\boldsymbol{\theta}(\mathbf{m}))$
11:    $\mathbf{m}\leftarrow$ Update$\left(\frac{\partial{\mathcal{L}_{upper}}}{\partial{\boldsymbol{\theta}(\mathbf{m})}},\mathbf{m},\boldsymbol{\theta}_{d},\eta_{2}\right)$     // Refer to Eq. (8)
@@ 4 Experiments
In our experiments, we use three open-source diffusion models, SD v1.4 [8], v1.5 [9] and v2.1 [10], to evaluate three domains of illegal model adaptations: 1) forging public figures’ portraits [22; 24], 2) duplicating copyrighted artworks [26] and 3) generating explicit content [25].
Datasets: For each domain, we use datasets as listed below, and random select different data classes as illegal and legal classes. We use 50% of samples in the selected classes for mask learning and model training, and the other samples for testing. More details about datasets are in Appendix C.
  • Portraits of public figures: We use a self-collected dataset, namely Famous-Figures-25 (FF25), with 8,703 publicly available portraits of 25 public figures on the Web. Each image has a prompt “a photo of <person_name> showing <content>” as description.
  • Copyrighted artworks: We use a self-collected dataset, namely Artwork, which contains 1,134 publicly available artwork images and text captions on the Web, from five famous digital artists with unique art styles.
  • Explicit contents: We use the NSFW-caption dataset with 2,000 not-safe-for-work (NSFW) images and their captions [1] as the illegal class. We use the Modern-Logo-v4 [5] dataset, which contains 803 logo images labeled with informative text descriptions, as the legal class.
Baseline schemes: Our baselines include full fine-tuning (FT), random tensor freezing, and two competitive unlearning schemes, namely UCE [23 and IMMA [65]. Existing data poisoning methods [59; 62; 51] cannot be used because all data we use is publicly online and cannot be poisoned.
@@
  • UCE [23]: It uses unlearning to guide the learned knowledge about illegal classes in the pre-trained model to be irrelevant or more generic.
@@
  • IMMA [65]: It reinitializes the model weights so that it is hard for users to conduct effective fine-tuning on the reinitialized model, in both illegal and legal classes.
Measuring image quality: We used FID [28] and CLIP [27] scores to evaluate the quality of generated images. In addition, to better identify domain-specific details in generated images, we also adopted domain-specific image quality metrics, listed as below and described in detail in Appendix D. For each text prompt, the experiment results are averaged from 100 generated images with different random seeds.
  • Domain-specific feature extractors: Existing work [54] reported that FID and CLIP fail to measure the similarity between portraits of human subjects, and cannot reflect human perception in images. Hence, for human portraits and artworks, we apply specific feature extractors on real and generated images, and measure the quality of generated images as cosine distance between their feature vectors. For portraits, we use face feature extractors (FN-L, FN, VGG) in DeepFace [50]. For artworks, we use a pretrained CSD model [52]. Details are in Appendix D.1.
  • NudeNet: We used NudeNet [2] to decide the probability of whether the generated images contain explicit contents, as the image’s safety score. Details are in Appendix D.2.
  • Human Evaluation: To better capture human perception in generated images, we recruited 16 volunteers with diverse backgrounds to provide human evaluations on image quality. For each image, volunteers scored how the generated image is likely to depict the same subject as in the real image from 1 to 7, where 1 means “very unlikely” and 7 means “very likely”. Details are in Appendix D.3.
Metric FN-L($\downarrow$) FN($\downarrow$) VGG($\downarrow$) FID($\downarrow$) Human ($\downarrow$)
Pre-trained model 0.96 0.92 0.93 164.8 -
Full FT illegal 0.436 0.455 0.581 144.6 6.7
legal 0.436 0.455 0.581 144.6 6.7
UCE illegal 0.445 0.464 0.598 152.9 4.6
legal 0.442 0.465 0.583 151.4 5.4
IMMA illegal 0.467 0.493 0.624 148.8 5.1
legal 0.462 0.475 0.610 145.9 5.8
FG-10% illegal 0.441 0.451 0.603 148.0 4.9
legal 0.429 0.45 0.585 143.6 6.2
R-10% illegal 0.433 0.451 0.588 143.7 6.8
legal 0.431 0.457 0.582 144.0 6.8
FG-30% illegal 0.482 0.504 0.631 153.7 3.6
legal 0.449 0.478 0.590 146.7 6.0
R-30% illegal 0.429 0.456 0.590 145.0 5.9
legal 0.429 0.456 0.590 145.0 5.9
FG-50% illegal 0.530 0.638 0.647 155.5 2.1
legal 0.499 0.527 0.608 149.5 4.3
R-50% illegal 0.513 0.543 0.638 151.6 3.7
legal 0.512 0.522 0.632 153.2 3.7
Table 2: Mitigation power in 10 illegal classes and 10 legal classes from the FF25 dataset, where worse image quality indicates stronger mitigation power. FG-$\rho$means using FreezeAsGuard to freeze $\rho$tensors and R-$\rho$means random freezing.
4.1 Mitigating Forgery of Public Figures’ Portraits
We evaluate FreezeAsGuard in mitigating forgery of public figures’ portraits, using FF25 dataset and SD v1.5 model. 10 classes are randomly selected from FF25 as illegal and legal classes, respectively. As shown in Table 2, FreezeAsGuard can mitigate illegal model adaptation by 40% compared to Full FT. When $\rho$ varies from 10% to 50%, it also outperforms the unlearning schemes by 37%, because these schemes cannot prevent relearning knowledge in illegal classes with new training data. It also ensures better legal model adaptation. With $\rho$=30%, the impact on legal adaptation is $<$5%.
Figure 7: Examples of public figures’ portraits generated by FreezeAsGuard under different freezing ratios ($\rho$)
When the freezing ratio ($\rho$) increases, the difference between FreezeAsGuard and random freezing diminishes, and their mitigation powers also reach a similar level. This means that only a portion of tensors are important for adaptation in specific illegal classes. With a high freezing ratio, random freezing is more likely to freeze these important tensors. Meanwhile, it could also freeze tensors that are important to legal classes, resulting in low performance in legal model adaptations. Hence, as shown in Figure 7, when $\rho$=30%, the mitigation power is high enough that the generated images no longer resemble those in training data, and further increasing $\rho$ could largely affect legal model adaptation.
Figure 8: Examples of generated public figures’ portraits by FreezeAsGuard with $\rho$=30and other baseline methods
Based on these results, we empirically consider $\rho$=30as the optimal freezing ratio on SD v1.5 for the domain of public figures’ portraits. Figure 8 shows example images of baseline methods and FreezeAsGuard with $\rho$=30%. We can find that FreezeAsGuard effectively prevents the generated images from being recognized as the subjects in illegal classes. Meanwhile, the fine-tuned model can still generate detailed background content and subjects’ postures aligned with the prompt, indicating that the mitigation power is highly selective and focuses only on subjects’ faces. More examples of generated images are in Appendix F.1.
Metric CSD($\downarrow$) FID($\downarrow$) CLIP($\uparrow$) Human($\downarrow$)
Pre-trained model 0.841 323.8 - -
Full illegal 0.347 187.6 32.31 5.9
legal 0.365 194.0 32.19 5.4
UCE illegal 0.426 190.9 32.28 3.3
legal 0.381 195.1 32.17 3.1
IMMA illegal 0.396 190.8 32.61 4.6
legal 0.377 195 32.98 5.1
FG-30% illegal 0.373 190.6 32.37 5.7
legal 0.382 194.1 32.10 5.2
R-30% illegal 0.351 186.7 32.45 5.6
legal 0.363 194.1 32.56 5.1
FG-50% illegal 0.453 194.5 32.04 3.5
legal 0.40 195.3 32.49 3.9
R-50% illegal 0.383 189.7 32.21 5.3
legal 0.405 196.0 32.43 3.7
FG-70% illegal 0.511 195.7 31.96 1.7
legal 0.41 195.3 32.58 3.8
R-70% illegal 0.441 189.2 32.12 4.9
legal 0.454 196.4 32.15 4.2
FG-85% illegal 0.574 201.2 31.74 1.6
legal 0.526 214.8 31.91 2.1
R-85% illegal 0.565 197.6 32.08 2.8
legal 0.586 210.4 32.09 2.7
Table 3: Mitigation power in one illegal class and one legal class from the Artwork dataset, where worse image quality indicates stronger mitigation power. FG-$\rho$means using FreezeAsGuard to freeze $\rho$tensors and R-$\rho$means random freezing.
Figure 9: Examples of artwork images generated by FreezeAsGuard with different freezing ratios
4.2 Mitigating Duplication of Copyright Artworks
We evaluate the capability of FreezeAsGuard in mitigating the duplication of copyrighted artworks, using the Artwork dataset and SD v2.1 model. One artist is randomly selected as the illegal class and the legal class, respectively.
The results with different freezing ratios are shown in Table 3 and Figure 9. Unlike results in Section 4.1 where data classes exhibit only subtle differences in facial features, different artists’ artworks demonstrate markedly different styles. Hence, a higher freezing ratio is required for sufficient mitigation power. We empirically decide the optimal freezing ratio for the domain of artwork is 70%. When $\rho$=70%, FreezeAsGuard can provide 47% more mitigation power in illegal classes compared to full fine-tuning, and 30% more compared to unlearning schemes. Figure 10 further shows example images generated by FreezeAsGuard with $\rho$=70%, and more examples can be found in Appendix F.2.
Figure 10: Examples of generated artworks by FreezeAsGuard with $\rho$=70and other baseline methods
Method Illegal Legal
NudeNet($\uparrow$) FID($\downarrow$) CLIP($\uparrow$)
Pre-trained model 0.47 - -
Full FT 1.29 158.1 32.79
UCE 1.20 158.5 30.07
IMMA 1.17 162.0 28.71
FG-30% 1.27 159.5 32.50
R-30% 1.30 158.8 32.79
FG-50% 1.06 163.2 31.83
R-50% 1.20 160.6 30.43
FG-70% 0.87 166.1 31.56
R-70% 1.12 161.8 28.66
FG-85% 0.85 166.5 30.34
R-85% 0.93 164.6 30.81
Table 4: Mitigation power in illegal class (NSFW-caption dataset) and legal class (Modern-Logo-v4 dataset), where worse image quality (in FID or CLIP) or lower NudeNet score indicates stronger mitigation power. FG-$\rho$means using FreezeAsGuard to freeze $\rho$tensors and R-$\rho$means random freezing.
4.3 Mitigating Generation of Explicit Contents
To evaluate FreezeAsGuard’s mitigation of explicit contents, we designate the NSFW-caption dataset as illegal class, and the Modern-Logo-v4 dataset as legal class. Results in Table 4 and Figure 11 show that, with $\rho$=70%, FreezeAsGuard significantly reduces the model’s capability of generating explicit contents by up to 38% compared to unlearning schemes, while maintaining the model’s adaptability in legal class. More image examples are in Appendix F.3.
4.4 Scalability of Mitigation Power
To evaluate FreezeAsGuard’s scalability over multiple illegal classes, we randomly pick 2, 5 and 10 public figures in the FF25 dataset, and 1, 2 and 3 artists in the Artworks dataset, as illegal classes. As shown in Table 5 and 6, when the number of illegal classes increases, FreezeAsGuard can retain strong mitigation power in both cases, and continuously outperforms the unlearning schemes. Note that, with more illegal classes, the difference of mitigation power between FreezeAsGuard and random freezing is smaller, because more illegal classes correspond to more adaptation-critical tensors, and random freezing is more likely to cover them.
Figure 11: Examples of generated images with explicit contents by FreezeAsGuard with $\rho$=70and other baseline methods
Method 2 classes 5 classes 10 classes
illegal legal illegal legal illegal legal
Full FT 0.397 0.397 0.424 0.424 0.436 0.436
UCE 0.435 0.444 0.443 0.437 0.445 0.442
IMMA 0.412 0.428 0.461 0.463 0.467 0.462
FG-30% 0.467 0.426 0.474 0.458 0.482 0.449
Table 5: Mitigation power in the FF25 dataset, measured by the FN-L score, with different numbers of illegal classes.
Method 1 class 2 classes 3 classes
illegal legal illegal legal illegal legal
Full FT 0.348 0.356 0.415 0.411 0.434 0.458
UCE 0.426 0.381 0.538 0.521 0.552 0.574
IMMA 0.396 0.377 0.483 0.463 0.536 0.496
FG-70% 0.511 0.410 0.609 0.473 0.648 0.525
Table 6: Mitigation power in the Artwork dataset, measured by the CSD score, with different numbers of illegal classes
4.5 The Learned Selection of Frozen Tensors
In Figure 12 and 13, we visualized the learned binary masks of tensor freezing for different illegal classes on the FF-25 and Artwork datasets, respectively, with the SD v1.5 model. These results show that on both datasets, the tensors being frozen for different illegal classes largely vary, indicating that our mask learning method can properly capture the unique tensors that are critical to each class, hence ensuring scalability. Note that in practice, no matter how many illegal classes are involved, the total amount of frozen tensors will always be constrained by the freezing ratio ($\rho$). When more illegal classes are involved, our results show that FreezeAsGuard is capable of identifying the most critical set of tensors for mitigating the fine-tuned model’s representation power.
4.6 Mitigation Power with Different Models
As shown in Table 7, when applied to different SD models, FreezeAsGuard constantly outperforms baseline schemes. SD v1.4 and v1.5 are generally stronger than SD v2.1, and the gap between illegal and legal classes in FreezeAsGuard is slightly better for v1.4 and v1.5 models. We hypothesize that better pre-trained models have more modularized knowledge distribution over model parameters, and hence allow FreezeAsGuard to have less impact on legal classes.
Method SD 1.4 SD 1.5 SD 2.1
illegal legal illegal legal illegal legal
Full 0.435 0.435 0.436 0.436 0.439 0.439
UCE 0.447 0.442 0.445 0.442 0.445 0.441
IMMA 0.451 0.448 0.467 0.462 0.463 0.454
FG-30% 0.489 0.453 0.482 0.449 0.474 0.450
Table 7: Mitigation power in the FF25 dataset, measured by the FN-L score, with different diffusion models
4.7 Reduction of Computing Costs
One advantage of freezing tensors is that it reduces the computing costs of fine-tuning. As shown in Table 8, when fine-tuning the model on a A6000 GPU, by applying FreezeAsGuard’s selection of tensor freezing, users can save 22%-48% GPU memory and 13%-21% wall-clock computing time, compared to other baselines without freezing ($\rho$=0%). Such savings, hence, well motivate users to adopt the FreezeAsGuard’s tensor freezing in their fine-tuning practices.